OpenStudy (mustap):

Geometry Question. Will give medal to whoever can give me a good answer. Thanks Use the Distance Formula and the x-axis of the coordinate plane. Show why the distance between two points on a number line (the x-axis) is | a – b |, where a and b are the x-coordinates of the points.

OpenStudy (mustap):

If anyone can help me that would be helpful. I have a small idea of what to do, but this question seems to confuse me for some reason.

OpenStudy (mustap):

@Saisuke<3 do you think you could help? you've been watching for a while now

OpenStudy (anonymous):

OH sorry!!!! i looked at the question but idk how to help u! im so so so sorry!!!! :(

OpenStudy (mustap):

oh thats fine I was just expecting you to start writing any second. Thanks anyways

OpenStudy (mustap):

@jim_thompson5910 @iambatman @SolomonZelman @pooja195 would appreciate some help from any one of you guys

jimthompson5910 (jim_thompson5910):

EVERY point on the x axis is of the form (x,0) where x is some number examples: (3,0) (2,0) (9,0) (27,0) etc

jimthompson5910 (jim_thompson5910):

The distance formula is \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] so you plug in the two points (a,0) and (b,0) into this formula above

OpenStudy (mustap):

so would it be \[d=\sqrt{(b-a)^2(0-0)^2}\]

OpenStudy (mustap):

how should I write this as the final answer? I'm pretty sure there has to be more to it.

jimthompson5910 (jim_thompson5910):

\[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] \[\Large d = \sqrt{\left(a-b\right)^2+\left(0-0\right)^2}\] \[\Large d = \sqrt{\left(a-b\right)^2+\left(0\right)^2}\] \[\Large d = \sqrt{\left(a-b\right)^2+0}\] \[\Large d = \sqrt{\left(a-b\right)^2}\] \[\Large d = |a-b|\] Note: \(\Large |a-b| = |b-a|\) so order does not matter in this particular case

jimthompson5910 (jim_thompson5910):

you forgot the + in between the parenthesis

OpenStudy (mustap):

oh yeah I forgot the plus.

jimthompson5910 (jim_thompson5910):

btw this trick works for any horizontal or vertical line. It doesn't have to be the x axis

OpenStudy (mustap):

so I just have to solve for d in the formula?

jimthompson5910 (jim_thompson5910):

it says "Show why the distance between two points on a number line (the x-axis) is | a – b |, where a and b are the x-coordinates of the points." and the steps are given above. So that's all you need to do really

OpenStudy (mustap):

I thought the question was more complicated then that

OpenStudy (mustap):

okay thanks alot!

jimthompson5910 (jim_thompson5910):

np